OpenStudy (anonymous):

Realllyyy need help with this question... The Cool Company determines that the supply function for its basic air conditioning unit is S(p) = 30 + 0.006p3 and that its demand function is D(p) = 150 - 0.12p2, where p is the price. Determine the price for which the supply equals the demand.

OpenStudy (anonymous):

\[S_{p}=30 + 0.006p^{3}\] \[D_p=150 - 0.12p^2\] to find p when Supply equals demand, set Supply equal to demand: \[S_p=D_p\] \[30 + 0.006p^{3}=150 - 0.12p^2\] now you have one equation with one variable, solve p for the solution